# Retrieval Helps or Hurts? A Deeper Dive into the Efficacy of Retrieval Augmentation to Language Models

### v2

@@ 4.1 Setup
Retrievers. We consider four retrievers that include BM25 Robertson et al. (2009), Contriever Izacard et al. (2021), BGE Xiao et al. (2023), GenRed Yu et al. (2023a) and an Oracle which retrieves the correct supporting passage for a given QA pair. BM25 is a static term-based sparse retriever that doesn’t require training. Contriever is a unsupervised dense retriever pre-trained on a large corpora and fine-tuned on MS-MARCO Nguyen et al. (2016). BGE is a supervised dense retriever pretrained on a large corpora and then fine-tuned on various supervised datasets including QA datasets such as NQ Kwiatkowski et al. (2019) and HotpotQA Yang et al. (2018). Both the retrievers retrieve from the Wikipedia corpus that is employed to create our benchmark. GenRead relies on prompting models to generate context for augmentation. And lastly, to measure the effectiveness of models reasoning capabilities we use Oracle retriever that retrieves the correct supporting passage for a given QA pair. Please refer to Appendix C.3 for more details.
@@ 4.2 Analysis of Model’s Recall Ability
Figure 6: Overall accuracy of Vanilla LMs and LMs augmented with GenRead, BM25, Contriever, BGE, and supporting passages (Oracle). Error bars show the $95\%$ bootstrap confidence intervals with $1000$ samples. Accuracy of Vanilla and GenRead improves with larger model sizes. Augmentation with BM25, Contriever, and BGE enhances accuracy, especially for smaller models. However, the gap diminishes and, in some cases, becomes negative for larger models such as the GPT series. Supporting passages (Oracle) prove beneficial for all models.
@@ 4.3 When Do Retrievers Help
Our observations indicate that BM25, Contriever, and BGE notably enhance the accuracy of small and medium models by up to $49.2\%$, while GenRead appears to be more effective for larger models with up to $4.1\%$ improvement. This suggests that the ability to maintain a coherent chain-of-thought emerges more prominently when the model size is sufficiently large.
We observe a significant improvement in accuracy for all models, irrespective of their sizes, when augmented with Oracle. Even the smallest model, Flan-T5-small, achieves an accuracy of $78.1\%$ with Oracle passages. However, substantial differences are observed between performances with BM25/Contriever/BGE and Oracle, which can be attributed to retrieval errors. When augmented with incorrect passages, most models struggle to provide correct answers.
@@
For example, Flan-T5-small answers “Susanna Wesley” while the true answer is “Mary White”. Note that the context supports the true answer, but Flan-T5-small cannot answer correctly. Highlighted by the example, we observe that small models tend to extract an entity as an answer, which appears near from the words in questions, leading to lower accuracy compared to larger models.
@@ 4.4 Deep Dive into Errors
Figure 7 shows the accuracy of GPT-3.5 on questions with different subject-relation (S-R) counts. Notably, the accuracy of Vanilla and GenRead is considerable for popular questions but significantly declines when S-R counts drop below $10$. This drop is likely due to models memorizing popular facts. Conversely, RALMs utilizing BM25, Contriever, and BGE exhibit greater robustness on less popular questions. However, for popular questions, their performance is inferior compared to the Vanilla model without augmentation. Given that supporting passages (Oracle) consistently enhance the performance of models, we hypothesize that retrieval errors for popular questions negatively impact the performance of RALMs using BM25, Contriever, and BGE.
@@
To verify the above hypothesis, we closely examine the relationship between RALM performance and retrieval errors for questions of varying popularity. To estimate retriever performance, we compute passage accuracy, marking a passage as correct if any sub-string of the passage is an exact match for a gold answer. We report the passage accuracy of different retrieval models in Figure 8. This trend aligns closely with RALM performance, as illustrated in Figure 7. Specifically, agreement ratios between RALM performance and passage accuracy are $85.5\%,82.4\%$, $88.7\%$, and $93.1\%$ for BM25, Contriever, BGE, and Oracle, respectively. Upon closer examination, we find that although retriever accuracy is low for rare questions, the drop is less significant compared to Vanilla LMs (see the leftmost plots in Figures 7 and 8). This indicates that retrieval augmentation is still beneficial for rare facts compared to recall. In contrast, Vanilla LMs perform better than retrievers for popular questions.
Figure 9: Analysis on Vanilla LMs with BM25, Contriver, and BGE passage accuracy over S-R counts and S counts ($n=$ the number of questions in the group). In the top row, S-R counts are higher than the median. In the bottom row, they are less than or equal to the median. In the left column, S counts are less than or equal to the median and in the right column, they are higher than the median.
@@
Figure 9 shows the accuracy of various models along with BM25, Contriver, and BGE passage accuracy across the four groups. Note that we plot QA accuracy and passage accuracy since we use the same criteria for them. We observe that a supervised retriever BGE outperforms vanilla LMs in all groups. This is attributed to the fact that it was fine-tuned with QA datasets about Wikipedia. For the generalizability to broader domains, we discuss the results with an unsupervised retriever BM25 in the following. [^8] Similar results were observed for Contriever; hence, we focus on BM25 for simplicity.

### v3

@@ 4.1 Setup
Retrievers. We consider seven retrievers that include BM25 Robertson et al. (2009), Contriever Izacard et al. (2021), GTR-large, GTR-xl Ni et al. (2022), BGE [^7] BAAI/bge-large-en Xiao et al. (2023), GenRead Yu et al. (2023a) and Oracle. BM25 is a static term-based sparse retriever that doesn’t require training. Contriever is a unsupervised dense retriever pre-trained on a large corpora and fine-tuned on MS-MARCO Nguyen et al. (2016). GTR and BGE are supervised dense retrievers pretrained on a large corpora and then fine-tuned on various supervised datasets including NQ Kwiatkowski et al. (2019) and HotpotQA Yang et al. (2018). Both GTR and BGE retrieve from the Wikipedia corpus that is also employed to create our benchmark. GenRead generates relevant passages by prompting large language models. Oracle retriever always retrieves the correct supporting passage for a given QA pair. We include it to measure the reasoning capabilities of the models in isolation from retriever errors. Please refer to Appendix C.3 for more details.
@@ 4.2 Analysis of Model’s Recall Ability
Figure 6: Overall accuracy of Vanilla LMs and LMs augmented with GenRead, BM25, Contriever, GTR-large, GTR-xl, BGE, and supporting passages (Oracle). Error bars show the $95\%$ bootstrap confidence intervals with $1000$ samples. Accuracy of Vanilla and GenRead improves with larger model sizes. Augmentation with retrievers enhances accuracy, especially for smaller models. However, the gap diminishes and, in some cases, becomes negative for larger models such as the GPT series. Supporting passages (Oracle) prove beneficial for all models.
@@ 4.3 When Do Retrievers Help
Our observations indicate that BM25, Contriever, GTR, and BGE notably enhance the accuracy of small and medium models by up to $49.2\%$, while GenRead appears to be more effective for larger models with up to $4.1\%$ improvement. This suggests that the ability to maintain a coherent chain-of-thought emerges more prominently when the model size is sufficiently large.
We observe a significant improvement in accuracy for all models, irrespective of their sizes, when augmented with Oracle. Even the smallest model, Flan-T5-small, achieves an accuracy of $78.1\%$ with Oracle passages. However, substantial differences are observed between performances with BM25/Contriever/GTR/BGE and Oracle, which can be attributed to retrieval errors. When augmented with incorrect passages, most models struggle to provide correct answers.
@@
For example, Flan-T5-small answers “Mary White” while the true answer is “Susanna Wesley”. Note that the context supports the true answer, but Flan-T5-small cannot answer correctly. Highlighted by the example, we observe that small models tend to extract an entity as an answer, which appears near from the words in questions, leading to lower accuracy compared to larger models.
@@ 4.4 Deep Dive into Errors
Figure 7 shows the accuracy of GPT-3.5 on questions with different subject-relation (S-R) counts. Notably, the accuracy of Vanilla and GenRead is considerable for popular questions but significantly declines when S-R counts drop below $10$. This drop is likely due to models memorizing popular facts. Conversely, RALMs utilizing BM25, Contriever, GTR, and BGE exhibit greater robustness on less popular questions. However, for popular questions, their performance is inferior compared to the Vanilla model without augmentation. Given that supporting passages (Oracle) consistently enhance the performance of models, we hypothesize that retrieval errors for popular questions negatively impact the performance of RALMs using BM25, Contriever, GTR, and BGE.
@@
To verify the above hypothesis, we closely examine the relationship between RALM performance and retrieval errors for questions of varying popularity. To estimate retriever performance, we compute passage accuracy, marking a passage as correct if any sub-string of the passage is an exact match for a gold answer. We report the passage accuracy of different retrieval models in Figure 8. This trend aligns closely with RALM performance, as illustrated in Figure 7. Specifically, agreement ratios between RALM performance and passage accuracy are $85.5\%,82.4\%$, $88.5\%$, $88.4\%$, $88.7\%$, and $93.1\%$ for BM25, Contriever, GTR-large, GTR-xl, BGE, and Oracle, respectively. Upon closer examination, we find that although retriever accuracy is low for rare questions, the drop is less significant compared to Vanilla LMs (see the leftmost plots in Figures 7 and 8). This indicates that retrieval augmentation is still beneficial for rare facts compared to recall. In contrast, Vanilla LMs perform better than retrievers for popular questions.
Figure 9: Analysis on Vanilla LMs with BM25, Contriver, GTR, and BGE passage accuracy over S-R counts and S counts ($n=$ the number of questions in the group). In the top row, S-R counts are higher than the median. In the bottom row, they are less than or equal to the median. In the left column, S counts are less than or equal to the median and in the right column, they are higher than the median.
@@
Figure 9 shows the accuracy of various models along with BM25, Contriver, GTR, and BGE passage accuracy across the four groups. Note that we plot QA accuracy and passage accuracy since we use the same criteria for them. We observe that supervised retrievers, GTR and BGE, outperform vanilla LMs in all groups. This is attributed to the fact that it was fine-tuned with QA datasets about Wikipedia. For the generalizability to broader domains, we discuss the results with an unsupervised retriever BM25 in the following. [^9] Similar results were observed for Contriever; hence, we focus on BM25 for simplicity.
@@ Appendix B Prompts for Question Generation/Refinement and Experiments
Generate a background document from Wikipedia to answer the given question. {{ question }}"